Sol Summary – February 3rd

Sol 6

Summary Title: A Blustery Day on Mars
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander
Mission Status: Everything is fine, how are you?

Sol Activity Summary:
A two-EVA day today kept the crew busy all day. The morning EVA (Renee, Julia, and I) went North to discover the "Goblins" in Lith Canyon. The crew followed an ancient stream bed into the canyon, a familiar site for myself once we reached a few of the more memorable dips, with a grand opening from rolling red cliffs to an open field followed by a deep canyon drop area near what we hope to confirm was the correct location.

The afternoon EVA (Zac, Tat, and SJ) can be summarized by two words, windy bubbles. EVA double-oh-seven ventured by foot North and scrambled up the red cliffs near Sagan Road to reach the Hab Ridge Road. Performing a variety of bubbles experiments, the crew experience upper wind limits for their expedition. They walked south towards the Hab for a scenic view of our campus from above.

Meanwhile back in the Hab, I was in full science mode taking dust measurements in the airlock and uploading biometric data to Earth. The winds are expected to keep the Hab cool overnight leading into Super Science Sunday… I mean Super Bowl Sunday. The crew is planning to coordinate outreach activities to see if we can help raise awareness for human spaceflight and Mars-forward research.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
EVA 8 is being planned to take the crew back up on the Hab Ridge for sample collection and scenic views. The ERAU spacesuit mobility study is being planned for video capture in the science dome. IN the afternoon we will have our next installment of the human factors study. During the day the team will be trying to post awesome social media content for #SuperScienceSunday.

Have a better hashtag?

Let us know on Facebook: https://www.facebook.com/TeamISUonMars and Twitter: https://twitter.com/ISUonMars.

Anomalies in work:
Robotic observatory currently not functional. MDRS Astronomy lead working problem.

Weather:
A very windy day. Skies were mostly clear.

Crew Physical Status:
The crew is in the groove zone.

EVA:
One is planned for tomorrow.

Reports to file:
1. EVA #8 request
2. EVA #6 and #7 Report
3. Ops Report
4. Sol Summary
5. Journalist Report
6. Green Hab Report
7. Daily Photos

Support Requested:
Heater for science dome requested.
There are potentially two packages at Hollow Mountain that we would love to have picked up.
Water tank refills will be needed in the next few days.